Suppose that you have 155 mL of a buffer that is 0.4800 M in both propanoic acid (C2H5COOH) and its conjugate base (C2H5COO−). Calculate the maximum volume of 0.4500.450 M HCl that can be added to the buffer before its buffering capacity is lost.
Answer #1

Buffer capacity is said to be lost if the pH of the buffer changes by 1 unit. Given that conc. of buffer = 0.480 M volume of

Number of moles of C2H3COOH = M*V = 0.480 M * 155 mL = 74.4 mmol Number of moles of C,H,COO-= M*V = 0.480 M * 155 mL = 74.4 m